# Learning Agility Adaptation for Flight in Clutter

### v1

@@ VII-A1 Comparison of the proposed framework with constant agility baselines and ablation tests
EVA-planner tends to plan (sometimes overly) conservative velocities, which can be seen in Fig. 7, while the proposed approach makes better use of the vehicle’s maneuverability, at the cost of large changes in velocity, as seen in Fig. 7.
More importantly, since the hand-designed cost function of EVA-planner does not capture some aspects of the problem, such as the necessity of risk and perception awareness, an unsafe decision may be output by the planner. An example is illustrated in Fig. 8. In such a case, the agility policy learns to behave cautiously when the vehicle plans to turn its head into area out of its perception range, which is caused by occlusion of a corner and attitude angle changes. In contrast, EVA-planner plans an unsafe velocity due to its illusion that there are no obstacles nearby, which is caused by occlusion, yaw angle changes, and perception latency. However, such perception-aware behaviors does not always occur in the policy obtained using only the first stage of reward. An example is shown in Fig. 5, where acceleration at a corner causes the vehicle to collide.
Figure 9: Flight to avoid a wall-like obstacle with a hidden pillar. The planned trajectory (red curve), the executed trajectory (colored curve), and the local map (blue) are indicated in the bottom right corner of each figure. (a) Flight with constant agility ${v}^{\dagger}=2.5$m/s. (b) Flight with constant agility ${v}^{\dagger}=3.5$m/s (c) Flight with agility adaptation.
Figure 10: Flight through dense and large-sized obstacles. (a) Flight with constant agility ${v}^{\dagger}=2.5$m/s. (b) Flight with agility adaptation.
@@ VII-B Real-World Experiments
We deploy the policy on a micro drone with a size of $\times 17$cm×$\times 10$cm. The drone is equipped with an onboard RealSense D430 depth camera. Computation/inference is performed on an onboard Jetson Orin NX module. We use the NOKOV motion capture system and a VIO to obtain the state estimates respectively in indoor and outdoor scenarios. The parameter setup is aligned with that in section VI-A.
@@ VII-B1 Wall-like obstacles with an obstacle hidden behind
We first test the policy in a representative scenario where a wall-like obstacle blocked near half of the vehicle’s FOV, and only after the vehicle avoids the corner could it observe the obstacle immediately afterward.
In such a scenario, the policies set to constant agility levels, e.g., in Fig. 9, $v^{\dagger}=2.5$m/s and $v^{\dagger}_{t}=3.5$m/s, are able to plan a collision-free trajectory in most of the trials, which is due to the efficiency of the EGO-planner itself. However, the planned trajectory is far from being able to be perfectly tracked by the low-level controller since the trajectory requires sudden changes of motion, which leads to collisions. In this case, the trajectory optimization cost function of EGO-planner [17, Eq. (8)] (somehow) sacrifices dynamical feasibility to ensure a collision-free trajectory, but still fails to spare the vehicle.
In contrast, as shown in Fig. 9, after plugging the learned policy, the vehicle exhibits perception-aware behaviors, slowing down in the corner of the wall and recovering agility as more extents of the hidden obstacle is observed, until it eventually slows down to stop at the goal.
@@ VII-B2 Artificial Clutters
We evaluate the policy in an artificial scenario shown in Fig. 10. In this scenario, a dense clutter and a large-sized obstacle exist in the first and second half of the scene, respectively. The constant agility policies, as shown in the Fig. 10, fails in half of the trials due to planning time overruns, the occurrence of emergence stops (as explained in section V-C), and also collisions caused by the same reason as that described in the previous subsection. With the learned module, the policy reasonably exploits agility in open areas, and also behaviors cautiously among dense obstacles, as shown in Fig. 10, thus ensuring success in almost all trials.
@@ VII-B3 Natural Clutters
We also evaluate the policy in natural clutters as shown in Fig. LABEL:fig:head, where the highly uneven environment triggers rich patterns of vehicle’s behavior. Despite the particularly complex environment, our policy exhibits aggressive behavior in open areas (Fig. LABEL:fig:head, top left and bottom right), but stays cautious to ensure safety when crossing narrow gap (Fig. LABEL:fig:head, top right) or among obstacles (Fig. LABEL:fig:head, bottom left).
@@ VIII Conclusion
We proposed a hierarchical learning and planning framework for agility adaptive flight in cluttered environments. On the one hand, the hierarchical framework allows the system to obtain a strong overall performance thanks to the existing powerful model-based trajectory planner. On the other hand, by doing so, we complement the ’missing jigsaw puzzle’ in traditional trajectory planners, thus unlocking the potential of the algorithmic system while freeing humans from hand-tuning labor. We see this as a necessary step towards a self-closing loop for autonomous flight vehicles.

### v2

@@ VII-A1 Comparison of the proposed framework with constant speed constraint baselines and ablation tests
EVA-planner tends to plan (sometimes overly) conservative velocities, which can be seen in Fig. 6b, while the proposed approach makes better use of the vehicle’s maneuverability, at the cost of large changes in velocity, as seen in Fig. 6a.
More importantly, since the hand-designed cost function of EVA-planner does not capture some aspects of the problem, such as the necessity of perception awareness, an unsafe decision may be output by the planner. An example is illustrated in Fig. 7. In such a case, the outer-loop policy learns to behave cautiously when the vehicle plans to turn its head into area out of its perception range, which is caused by occlusion of a corner and attitude angle changes. In contrast, EVA-planner plans an unsafe velocity due to its illusion that there are no obstacles nearby, which is caused by occlusion, yaw angle changes, and perception latency. However, such perception-aware behaviors does not always occur in the policy obtained using only the first stage of reward. An example is shown in Fig. 4c, where acceleration at a corner causes the vehicle to collide.
Figure 8: Flight to avoid a wall-like obstacle with a hidden pillar. The planned trajectory (red curve), the executed trajectory (colorful curve), and the local map (blue) are indicated in the bottom right corner of each figure. (a) Flight with constant speed constraint ${v}^{\dagger}=2.5$m/s. (b) Flight with constant speed constraint ${v}^{\dagger}=3.5$m/s (c) Flight with speed adaptation.
Figure 9: Flight through dense and large-sized obstacles. (a) Flight with constant speed constraint ${v}^{\dagger}=2.5$m/s. (b) Flight with speed adaptation.
@@ VII-B Real-World Experiments
We deploy the policy on a micro drone with a size of $\times 17$cm×$\times 10$cm. The drone is equipped with an onboard RealSense D430 depth camera. Computation is performed on an onboard Jetson Orin NX module. We use the NOKOV motion capture system and a VIO to obtain the state estimates in indoor and outdoor scenarios, respectively. The parameter setup is aligned with that in section VI-A.
@@ VII-B1 Wall-like obstacles with an obstacle hidden behind
We first test the policy in a representative scenario where a wall-like obstacle blocks near half of the vehicle’s FOV, and only after the vehicle avoids the corner could it observe the obstacle immediately afterward.
We find that, in such a scenario, the policies under constant speed constraints, e.g., $v^{\dagger}=2.5$m/s and $v^{\dagger}_{t}=3.5$m/s in Fig. 8, can plan a collision-free trajectory in most of the trials. However, the mapping and planning pipeline may exhibit a significant perception latency and the planner may generate a trajectory that is not dynamically feasible. As a result, the planned trajectory, which requires sudden changes of motion, can be far from being able to be perfectly tracked by the low-level controller. We note that, typically, the trajectory is dynamically feasible and thus can be well-tracked, while in some emergency cases, the trajectory optimization method of EGO-planner [12, Eq. (8)], which uses penalties to surrogate constraints for computational efficiency, can sacrifice dynamical feasibility to ensure a collision-free trajectory.
In contrast, as shown in Fig. 8c, after plugging the learned policy, the vehicle exhibits perception-aware behaviors, slowing down at the corner of the wall and recovering aggressiveness as more extents of the hidden obstacle is observed, until it eventually slows down to stop at the goal.
@@ VII-B2 Artificial Clutters
We evaluate the policy in an artificial scenario shown in Fig. 9. In this scenario, a dense clutter and a large-sized obstacle exist in the first and second half of the scene, respectively. The constant speed constraint policies, as shown in the Fig. 9a, fails in half of the trials due to planning time overruns, the occurrence of emergence stops (as explained in section V-C), and also collisions caused by the same reason as that described in the previous subsection. With the learned module, the policy reasonably exhibits aggressiveness in open areas, and also behaviors cautiously among dense obstacles, as shown in Fig. 9b, thus ensuring success in almost all trials.
@@ VII-B3 Natural Clutters
We also evaluate the policy in natural clutters as snapshotted in Fig. LABEL:fig:head, where the highly uneven environment triggers rich patterns of vehicle’s behavior. Despite the particularly complex environment, our policy exhibits aggressive behavior in (relatively) open areas (Fig. LABEL:fig:head, top left and bottom right), but stays cautious to ensure safety when crossing narrow gap (Fig. LABEL:fig:head, top right) or among obstacles (Fig. LABEL:fig:head, bottom left).
@@ VIII Conclusion
We propose a hierarchical learning and planning framework for aggressiveness-adaptive flight in cluttered environments. On the one hand, the hierarchical framework allows the system to obtain a strong overall performance thanks to the existing powerful model-based trajectory planner. On the other hand, by doing so, we complement the ’missing jigsaw puzzle’ in traditional trajectory planners, thus unlocking the potential of the system while freeing humans from hand-tuning labor.
The main limitation of the current system is that the spatial distribution of the vehicle is determined by the trajectory planner itself, and thus in some adversarial environments the sub-optimality of the design (e.g., choice of topology) in the planner backbone may largely limit the allowed aggressiveness in the environment. To achieve versatile aggressive flight in any environment, one solution is to give greater dominance to the learned policy while still taking advantage of safety guarantees and generalizability rendered by the traditional planner. However, when the search space is expanded, learning can take place less efficiently especially when embedded in a model-based planner where the depth rendering and simulating the system are both time-consuming. Moreover, the specialized reward scheme used in this work may not be generalized when considering a different auxiliary action space. Therefore, both scalable training environment and a unified method for effective learning are highly desired.